OpenStudy (anonymous):

k^2+4k=96 trinomials

OpenStudy (geometry_hater):

you need to set up your problem the right way.First off you need to subtract 95 So it looks like this K^2 +4K -96 = 0 You must factor

OpenStudy (anonymous):

but that the thing im so confused where do i start first like how do i do it

OpenStudy (anonymous):

You could use the COMPLETING THE SQUARE method k^2+4k=96 add 2^2 to both sides k^2 + 4k + 2^2 = 96 + 2^2 LHs becomes a perfect square (k + 2)^2, so we get (k + 2)^2 = 96 + 4 taking square root of both sides k + 2 = +/- sqrt(100) k = -2 +/- 10 so k = -2 - 10 = -12 or k = -2 + 10 = 8

OpenStudy (anonymous):

For factorising, first rearrange the equation to read K^2 +4K -96 = 0 Now find factor pairs of -96 (i.e. constant at the end) and select the pair whose product is - 96 sum is + 4 (coefficient of the x-term ) -1 x 96 or 1 x -96 -2 x 48 or 2 x -48 -3 x 32 or 3 x -32 -4 x 24 or 4 x -24 -6 x 16 or 6 x -16 -8 x 12 or 8 x -12 Out of these we find the pair -8 x 12 = -96 and -8 + 12 = +4 so we split the middle term as follows k^2 -8k + 12k - 96 = 0 k(k - 8) + 12(k - 8) = 0 (k - 8)(k + 12) = 0 using zero product property if k - 8 = 0, then k = 8 if k + 12 = 0, then k = -12